# Usefulness of Quantum Entanglement for Enhancing Precision in Frequency Estimation

### v1

@@ II Identifiability limits: the usefulness of GHZ states
Comparison of bounds (7) and (4) shows that the use of initial maximally entangled states is disadvantageous when compared with the use initial product states of the probes, if the only resource taken into account is the total number $\nu N=\nu N_{\textrm{G}HZ}$ of probes. Indeed, the maximal reachable precision in the estimation of $\omega$ is worse by a factor $\sqrt{N}$
@@
Figure 1: Adaptive-time frequency estimation. The first step consist of a measurement over a phase codified in a quantum state. If the error in the measurement is smaller than the step goal, the measurement time is incremented, the goal is updated and the process is repeated, in this case, the quantum information for the next measurment is larger. If the error in the measurement is equal or larger than the step goal, the measurement is repeated without incrementing the measurement time. This procedure is repeated until the global error goal is achieved.
@@ III.1 Optimal measurements and asymptotic consistent estimators
We determine now the estimation error of ATFE. For a given confidence level $C_{l}\in\left[0,1\right]$, in each adaptive step, we have two contributions to the estimation error: i) when the confidence interval $CI=(\widehat{\omega}-E,\widehat{\omega}+E)$, where $E$ is the marginal error and $\widehat{\omega}$ is the estimation, contains the value of the parameter and ii) when it does not. When the confidence interval contains the value of the parameter, the MLE from AQSE converges in distribution to a normal distribution with mean at $\tilde{\omega}$ and variance equal to the inverse of the Fisher information [35, 21]. On the other hand, if the confidence interval does not contain the parameter, AQSE produces an error larger than $E$ [22].
@@
Figure 3: Lower bound in the Holevo variance for the maximum likelihood estimator produced by ATFE. Eq. (32) is plotted for the confidence levels $C_{l}=0.95$ (yellow line), $C_{l}=0.99$ (blue line), and $C_{l}=0.999$ (green line). The cross markers in the curves indicate the adaptive step when the time is increased in the method.
@@ Acknowledgments
We thank Laboratorio Universitario de Cómputo de Alto Rendimiento (LUCAR) of IIMAS-UNAM for their service on information processing. MR acknowledges funding from the National Science Foundation (NSF) grants FRHTP # PHY-2116246, and # 2016244 - QLCI-CI. PB-B gratefully acknowledges support from the PAPIIT-DGAPA grant IG101324.

### v2

@@ II Identifiability limits: the usefulness of GHZ states
Comparison of bounds (7) and (4) shows that the use of initial maximally entangled states is disadvantageous when compared with the use of initial product states of the probes, if the only resource taken into account is the total number $\nu N=\nu N_{\textrm{G}HZ}$ of probes. Indeed, the maximal reachable precision in the estimation of $\omega$ is worse by a factor $\sqrt{N}$
@@
Figure 1: Adaptive-time frequency estimation. The procedure begins with a measurement of a phase encoded in a quantum state. If the measurement error is smaller than the desired threshold (local error) for the current step, the measurement time is increased, the threshold is recalculated, and the process is repeated. In this case, the quantum Fisher information available for the subsequent measurement increases. Conversely, if the measurement error is equal to or exceeds the current threshold, the measurement is repeated without adjusting the measurement time. This procedure is repeated until the desired overall error (global error) is achieved.
@@ III.1 Optimal measurements and asymptotic consistent estimators
We determine now the estimation error of ATFE. For a given confidence level $C_{l}\in\left[0,1\right]$, in each adaptive step, we have two contributions to the estimation error: i) when the confidence interval $\widehat{\textrm{CI}}=(\widehat{\omega}-E,\widehat{\omega}+E)$, where $E$ is the marginal error and $\widehat{\omega}$ is the estimation, contains the value of the parameter and ii) when it does not. When the confidence interval contains the value of the parameter, the MLE from AQSE converges in distribution to a normal distribution with mean at $\tilde{\omega}$ and variance equal to the inverse of the Fisher information [35, 21]. On the other hand, if the confidence interval does not contain the parameter, AQSE produces an error larger than $E$ [22].
@@
Figure 3: Lower bound in the Holevo variance for the MLE produced by ATFE. Eq. (32) is plotted for the confidence levels $C_{l}=0.95$ (yellow line), $C_{l}=0.99$ (blue line), and $C_{l}=0.999$ (green line). The cross markers in the curves indicate the adaptive step when the time is increased in the method.
@@
Extending Ramsey phase identifiability
For product states, when individual qubits can be independently controlled and measured, the initial phase interval where the Ramsey phase $\phi$ is identifiable can be extended. This larger interval can be used to extend the evolution time, $\tilde{t}$, at each adaptive step in the ATFE protocol. One such strategy is the dual-quadrature measurement introduced in Ref. [40], which, by controlling and measuring two qubits, can expand the phase interval from $[-\pi/2,\pi/2]$ to $[-\pi,\pi]$. Furthermore, larger extensions of the phase interval—and consequently longer evolution times, $\tilde{t}$—can be achieved by using $N$ atoms and measuring them over different timescales. This protocol allows the evolution time to increase by factors of $2,\ldots,2^{N/2-1}$m with $N$ even.
Using this technique, the maximum measurement time $t_{\textrm{prod}}$ in the ATFE protocol can be extended from $\pi/(2\Delta\Omega)$ to $2^{N/2}\pi/(2\Delta\Omega)$, for $N$ even. This extension does not alter the qualitative behavior of the error predicted by our method; however, by allowing for larger time steps, it reduces the error more rapidly. The trade-off is the additional complexity required to individually control and measure each qubit. In Figure 5, we show how the ATFE protocol improves when the dual-quadrature measurement is implemented for different numbers of atoms.
Figure 5: Variance of the MLE, $\text{Var}(\hat{\omega}_{\text{MLE}})\cdot\nu$, using different number of atoms with the dual-quadrature measurement described in [40] combined with the ATFE protocol. The variance is shown as a function of the total time $T$ (left) and the number of measurements $\nu$ (right) for different particle numbers, $N={1,2,4,8}$. The cases for even $N$ correspond to the strategy of dual-quadrature measurements, while the curve for $N=1$ represents the ATFE with standard Ramsey spectroscopy, used here as a reference threshold. Both figures demonstrate that the ATFE protocol reduces the error compared to the case without adaptive measurements. As $T$ or $\nu$ increases, the variance decreases, reaching a minimum defined by the confidence interval and the initial time $\tilde{t}_{1}=1/2$ for $N=1$ and $\tilde{t}_{1}=2^{N/2-1}$ for even $N$. The qualitative behavior of the variance remains consistent for different numbers of qubits in both figures. The gray solid lines indicate the estimation error predicted by the QCRB for Ramsey spectroscopy when the ATFE protocol is not implemented.
@@ Acknowledgments
We thank Laboratorio Universitario de Cómputo de Alto Rendimiento (LUCAR) of IIMAS-UNAM for their service on information processing. M.R.-G. acknowledges funding from the National Science Foundation (NSF) Grants FRHTP No. PHY-2116246 and No. QLCI-2016244. P.B.-B. gratefully acknowledges support from the PAPIIT-DGAPA Grant No. IG101324. R.L.M.F. acknowledges the support of the John Templeton Foundation (Grant No. 62424).